the sqare root of 2x minus 1 =2
Answer by mananth(16946)

sqrt(2x-1)=2
square both sides
2x-1=4
add1 to both sides
2x-1+1=4+1
2x=5
divide by 2
2x/2 = 5/2
x=5/2